if the sum of the edges of a cube is 144 cm, what is the volume of the same cube

Found 2 solutions by algebrahouse.com, Alan3354:
Answer by algebrahouse.com(1659)

"if the sum of the edges of a cube is 144 cm, what is the volume of the same cube"

x + x + x = 144 {sum of edges is 144}
3x = 144 {combined like terms}
x = 28 {divided both sides by 3}

V = x³ {volume of a cube is one side cubed}
= 28³
= 28(28)(28)
= 21,952 cm³
www.algebrahouse.com

Answer by Alan3354(69443)

if the sum of the edges of a cube is 144 cm, what is the volume of the same cube
---------------
A cube has 12 edges.
144/12 = 12 cm
Volume = 12^3 = 1728 cc